# Relative Entropy Methods for Calculating Committors

### v1

@@ Appendix A Proofs of Results in Section 3
Proof of Theorem 1.
Since we assume the Novikov condition, by Lemma 5,
@@
is an $\mathcal{F}_{t}$-martingale under $\mathds{Q}$ with $\mathds{Q}[M_{t}]=1$ for $t\geq 0$. It follows that
$Z_{t}$ $=\frac{\mu^{-1}m(Y_{0})}{\nu^{-1}\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}M_{t}$
is a martingale with $\mathds{Q}[Z_{t}]=1$: For $0\leq s\leq t$, we have
$\mathds{Q}[Z_{t}|\mathcal{F}_{s}]$ $=\mathds{Q}\left[\left.\frac{\mu^{-1}m(Y_{0})}{\nu^{-1}\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}M_{t}\right|\mathcal{F}_{s}\right]$
$=\frac{\mu^{-1}m(Y_{0})}{\nu^{-1}\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}\mathds{Q}[M_{t}|\mathcal{F}_{s}]$
$=\frac{\mu^{-1}m(Y_{0})}{\nu^{-1}\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}M_{s}.$
@@
$\mathds{Q}[Z_{t}]$ $=\mathds{Q}\left[\mathds{Q}\left[\left.\frac{\mu^{-1}m(Y_{0})}{\nu^{-1}\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}M_{t}\right|\mathcal{F}_{0}\right]\right]$
$=\mathds{Q}\left[\frac{\mu^{-1}m(Y_{0})}{\nu^{-1}\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}\mathds{Q}[M_{t}|\mathcal{F}_{0}]\right]$
$=\mathds{Q}\left[\frac{\mu^{-1}m(Y_{0})}{\nu^{-1}\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}M_{0}\right]$
@@
since $M_{0}=1$ and $Y_{0}\sim\nu^{-1}\lvert\nabla q(x)\rvert\exp(-U(x)/\varepsilon)$ under $\mathds{Q}$.
We now show that $Z_{\tau}$ is the density of $\mathds{P}$ with respect to $\mathds{Q}$ restricted to the stopping time $\sigma$-algebra $\mathcal{F}_{\tau}$. Since $Z_{t}$ is a martingale, the optional stopping theorem [14, Chapter 2, Theorem 3.2] implies that for any fixed $t>0$,
$Z_{t\wedge\tau}=\mathds{Q}[Z_{t}|\mathcal{F}_{t\wedge\tau}].$
@@
$=\lim_{N\rightarrow\infty}\int_{A_{N}}\mathds{Q}[Z_{N}(\omega)|\mathcal{F}_{\tau(\omega)\wedge N}]\mathds{Q}(\mathrm{d}\omega)$
@@
and so $Z_{\tau}$ is the density of $\mathds{P}$ with respect to $\mathds{Q}$ on $\mathcal{F}_{\tau}$. ∎
Appendix B Proofs of Results in Section 4.1
First, we prove Lemma 7 which provides an alternative expression for the improper integral in our change of measure formula.
Proof of Lemma 7.
By Lemma 2, $w$ is smooth. Therefore, by Ito’s formula, under $\mathds{P}$, we have
$dw(Y_{t})=Lw(Y_{t})\,\mathrm{d}t+2\epsilon\nabla\log\tilde{q}(Y_{t})\cdot\nabla w(Y_{t})\,\mathrm{d}t+\sqrt{2\epsilon}\nabla w(Y_{t})\cdot\mathrm{d}W_{t},\$ (23)
where $W_{t}$ is the $\mathds{P}$-Brownian motion defined in Theorem 1 Now let
@@
By a calculation similar to the proof of Lemma 4, we have
@@
and the result follows. ∎
We use Lemma 7 to show that our estimator of relative entropy differences has finite variance, at least under certain conditions on $\tilde{q}$.
Lemma 8.
Assume that $\mathds{P}[\tau^{2}]<\infty$, that for some $C>0$,
$\left\lvert\log\left(\frac{\tilde{q}(x)}{q(x)}\right)\right\rvert\leq C\text{ and }\left\lvert\nabla\log\left(\frac{\tilde{q}(x)}{q(x)}\right)\right\rvert\leq C$
@@
$\mathds{P}\left[\left(\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right)^{2}\right]<\infty$
even though $\frac{L\tilde{q}}{\tilde{q}}$ may be singular on $\partial A$. Taking $T=q$ in Lemma 7 yields
@@
where $r=\tilde{q}/q$ and $W_{s}$ is the $\mathds{P}$-Brownian motion in Theorem 1. Therefore,
$\mathds{P}\left[\left(\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right)^{2}\right]^{\frac{1}{2}}$ $\leq\mathds{P}\left[\left(\log\frac{r(Y_{\tau})}{r(Y_{0})}\right)^{2}\right]^{\frac{1}{2}}+\varepsilon\mathds{P}\left[\left(\int_{0}^{\tau}\lvert\nabla\log r(Y_{s})\rvert^{2}\,\mathrm{d}s\right)^{2}\right]^{\frac{1}{2}}$
$\qquad\qquad\qquad+\sqrt{2\varepsilon}\mathds{P}\left[\left(\int_{0}^{\tau}\nabla\log r(Y_{s})\cdot\mathrm{d}W_{s}\right)^{2}\right]^{\frac{1}{2}}$
$\leq 2C+\varepsilon C\mathds{P}[\tau^{2}]+\sqrt{2\varepsilon}\mathds{P}\left[\left(\int_{0}^{\tau}\nabla\log r(Y_{s})\cdot\mathrm{d}W_{s}\right)^{2}\right]^{\frac{1}{2}}.$ (24)
@@
$\mathds{P}[(S_{\tau\wedge n^{\prime}}-S_{\tau\wedge n})^{2}]$ $=\mathds{P}\left[\left(\int_{\tau\wedge n}^{\tau\wedge n^{\prime}}\nabla\log r(Y_{s})\cdot\mathrm{d}W_{s}\right)^{2}\right]$
$=\mathds{P}\left[\left(\int_{0}^{n^{\prime}}\mathds{1}_{\tau\wedge n\leq s<\tau\wedge n^{\prime}}\nabla\log r(Y_{s})\cdot\mathrm{d}W_{s}\right)^{2}\right]$
$=\mathds{P}\left[\int_{0}^{n^{\prime}}\mathds{1}_{\tau\wedge n\leq s<\tau\wedge n^{\prime}}\lvert\nabla\log r(Y_{s})\rvert^{2}\,\mathrm{d}s\right]$
$\leq\mathds{P}\left[\int_{0}^{n^{\prime}}\mathds{1}_{\tau\geq n}\lvert\nabla\log r(Y_{s})\rvert^{2}\,\mathrm{d}s\right]$
@@
$\mathds{P}[S_{\tau}^{2}]$ $=\lim_{n\rightarrow\infty}\mathds{P}[S_{\tau\wedge n}^{2}]$
$=\lim_{n\rightarrow\infty}\mathds{P}\left[\left(\int_{0}^{n}\mathds{1}_{s\leq\tau\wedge n}\nabla\log r(Y_{s})\cdot\mathrm{d}W_{s}\right)^{2}\right]$
$=\lim_{n\rightarrow\infty}\mathds{P}\left[\int_{0}^{n}\mathds{1}_{s\leq\tau\wedge n}\lvert\nabla\log r(Y_{s})\rvert^{2}\,\mathrm{d}s\right]$
$\leq\lim_{n\rightarrow\infty}C^{2}\mathds{P}[\tau\wedge n]$
$=C^{2}\mathds{P}[\tau].$
@@
We leave the remaining details to the reader. ∎
We now prove Theorem 2 on the differentiability of $D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ in $\theta$.
Proof of Theorem 2.
@@
Moreover, since we assume $\mu_{\theta}\geq\frac{1}{M}>0$, and since $q_{\theta}$ is differentiable, $\log\mu_{\theta}$ is differentiable. Therefore,
$\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}$ $=\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\left(\frac{\nabla_{\theta}q_{\theta}(X_{\tau})}{q_{\theta}(X_{\tau})}-\frac{\nabla_{\theta}\mu_{\theta}}{\mu_{\theta}}-\int_{0}^{\tau}\nabla_{\theta}\frac{Lq_{\theta}}{q_{\theta}}(X_{s})\,\mathrm{d}s\right).$ (25)
@@
$D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})=\mathds{Q}\left[g\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)\right].$
@@
$\nabla_{\theta}D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})=\nabla_{\theta}\mathds{Q}\left[g\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)\right]=\mathds{Q}\left[g^{\prime}\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right].$ (26)
Assumption 7 does not imply that $\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}$ is bounded uniformly in $\theta$ by a function in $L^{1}(\mathds{Q})$, so textbook results related to differentiating under the integral do not apply. To verify (26), we return to the definition of the derivative. Given any $h,\theta\in\mathbb{R}^{k}$, by the mean value theorem, for some $s:\Omega\rightarrow[0,1]$, we have
$\frac{1}{\lVert h\rVert}\mathds{Q}$ $\left[g\left(\frac{\mathrm{d}\mathds{P}_{\theta+h}}{\mathrm{d}\mathds{Q}}\right)-g\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)-g^{\prime}\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\cdot h\right]$
$=\mathds{Q}\left[g^{\prime}\left(\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\right)\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\cdot\frac{h}{\lVert h\rVert}-g^{\prime}\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\cdot\frac{h}{\lVert h\rVert}\right]$
=[$=\mathds{Q}\left[\left\{g^{\prime}\left(\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\right)-g^{\prime}\left(\frac{\mathrm{d}\mathds{P}_{\theta+h}}{\mathrm{d}\mathds{Q}}\right)\right\}\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\cdot\frac{h}{\lVert h\rVert}\right.$
@@
$=\mathds{P}_{\theta}\left[\left\{\log\left(\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\right)-\log\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)\right\}\nabla_{\theta}\log\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\cdot\frac{h}{\lVert h\rVert}\right.$
+$\qquad\qquad+\left\{\log\left(\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\right)+1\right\}\nabla_{\theta}\log\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\cdot\frac{h}{\lVert h\rVert}\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\frac{\mathrm{d}\mathds{Q}}{\mathrm{d}\mathds{P}_{\theta}}$
@@
$=:\mathds{P}_{\theta}[R(\theta;h)].$
As $h\rightarrow 0$, the integrand $R(\theta;h)$ in the expectation with respect to $\mathds{P}_{\theta}$ above converges pointwise to zero. We now show that $R(\theta;h)$ is dominated in $L^{1}(\mathds{P}_{\theta})$, uniformly in $h$ for sufficiently small $h$. If so, $\lim_{h\rightarrow 0}\mathds{P}_{\theta}[R(\theta;h)]=0$, which verifies differentiability of $D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ and (26). By Assumption 7, we have
$\left\lvert\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\frac{\mathrm{d}\mathds{Q}}{\mathrm{d}\mathds{P}_{\theta}}\right\rvert$ $\leq\frac{\mu_{\theta}}{\mu_{\theta+sh}}\exp\left(\int_{0}^{\tau}\left\lvert\frac{Lq_{\theta}}{q_{\theta}}(Y_{t})-\frac{Lq_{\theta+sh}}{q_{\theta+sh}}(Y_{t})\right\rvert\,\mathrm{d}t\right)$
@@
$\left\lvert\log\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)\right\rvert$ $\leq\left\lvert\log\left(\frac{\mu_{\theta}}{\nu}\right)\right\rvert+\int_{0}^{\tau}\left\lvert\frac{Lq_{\theta}}{q_{\theta}}(Y_{t})\right\rvert\,\mathrm{d}t$
$\leq\lvert\log\nu\rvert+\lvert\log M\rvert+C\tau,$
@@
$\left\lvert\nabla_{\theta}\log\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right\rvert$ $\leq\left\lvert\frac{\nabla_{\theta}m_{\theta}}{m_{\theta}}\right\rvert+\int_{0}^{\tau}\left\lvert\nabla_{\theta}\frac{Lq_{\theta}}{q_{\theta}}(Y_{t})\right\rvert\,\mathrm{d}t$
@@
$\lvert R(\theta;h)\rvert$ $\leq 2(\lvert\log\nu\rvert+\lvert\log M\rvert+C\tau)(M^{2}+C\tau)$
$\qquad+(1+\lvert\log\nu\rvert+\lvert\log M\rvert+C\tau)(M^{2}+C\tau)$
$\qquad+(1+\lvert\log\nu\rvert+\lvert\log M\rvert+C\tau)(M^{2}+C\tau)M^{2}\exp\left(\frac{\gamma(\theta)}{2}\tau\right).$
Under Assumption 7, the random variable on the right-hand-side above is in $L^{1}(\mathds{P}_{\theta})$, which concludes the proof of differentiability.

### v2

@@ Appendix A Proofs of Results in Section 3
Proof A.1 (Proof of Theorem 3.11).
Since we assume the Novikov condition, by Lemma 3.8,
@@
is an $\mathcal{F}_{t}$-martingale under $\mathds{Q}$ with $\mathds{E}_{\mathds{Q}}[M_{t}]=1$ for $t\geq 0$. It follows that
$Z_{t}$ $=\frac{\eta^{-1}m(Y_{0})}{\zeta^{-1}\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}M_{t}$
is a martingale with $\mathds{E}_{\mathds{Q}}[Z_{t}]=1$: For $0\leq s\leq t$, we have
$\mathds{E}_{\mathds{Q}}[Z_{t}|\mathcal{F}_{s}]$ $=\mathds{E}_{\mathds{Q}}\left[\left.\frac{\eta^{-1}m(Y_{0})}{\zeta^{-1}\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}M_{t}\right|\mathcal{F}_{s}\right]$
$=\frac{\eta^{-1}m(Y_{0})}{\zeta^{-1}\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}\mathds{E}_{\mathds{Q}}[M_{t}|\mathcal{F}_{s}]$
$=\frac{\eta^{-1}m(Y_{0})}{\zeta^{-1}\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}M_{s}.$
@@
$\mathds{E}_{\mathds{Q}}[Z_{t}]$ $=\mathds{E}_{\mathds{Q}}\left[\mathds{E}_{\mathds{Q}}\left[\left.\frac{\eta^{-1}m(Y_{0})}{\zeta^{-1}\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}M_{t}\right|\mathcal{F}_{0}\right]\right]$
$=\mathds{E}_{\mathds{Q}}\left[\frac{\eta^{-1}m(Y_{0})}{\zeta^{-1}\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}\mathds{E}_{\mathds{Q}}[M_{t}|\mathcal{F}_{0}]\right]$
$=\mathds{E}_{\mathds{Q}}\left[\frac{\eta^{-1}m(Y_{0})}{\zeta^{-1}\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}M_{0}\right]$
@@
since $M_{0}=1$ and $Y_{0}\sim\zeta^{-1}\lvert\nabla q(x)\rvert\exp(-U(x)/\varepsilon)$ under $\mathds{Q}$.
We now show that $Z_{\tau}$ is the density of $\mathds{P}$ with respect to $\mathds{Q}$ restricted to the stopping time $\sigma$-algebra $\mathcal{F}_{\tau}$. Since $Z_{t}$ is a martingale, the optional stopping theorem [26, Chapter 2, Theorem 3.2] implies that for any fixed $t>0$,
$Z_{t\wedge\tau}=\mathds{E}_{\mathds{Q}}[Z_{t}|\mathcal{F}_{t\wedge\tau}].$
@@
$=\lim_{N\rightarrow\infty}\int_{A_{N}}\mathds{E}_{\mathds{Q}}[Z_{N}(\omega)|\mathcal{F}_{\tau(\omega)\wedge N}]\mathds{Q}(\mathrm{d}\omega)$
@@
and so $Z_{\tau}$ is the density of $\mathds{P}$ with respect to $\mathds{Q}$ on $\mathcal{F}_{\tau}$.
Appendix B Proofs of Results in Section 4
First, we prove Lemma 5.1 which provides an alternative expression for the improper integral in our change of measure formula.
Proof B.1 (Proof of Lemma 5.1).
By Lemma 3.1, $w$ is smooth. Therefore, by Ito’s formula, under $\mathds{P}$, we have
(33) $dw(Y_{t})=Lw(Y_{t})\,\mathrm{d}t+2\epsilon\nabla\log\tilde{q}(Y_{t})\cdot\nabla w(Y_{t})\,\mathrm{d}t+\sqrt{2\epsilon}\nabla w(Y_{t})\cdot\mathrm{d}W_{t},\\$
where $W_{t}$ is the $\mathds{P}$-Brownian motion defined in Theorem 3.11 Now let
@@ Appendix B Proofs of Results in Section 4
By a calculation similar to the proof of Lemma 3.6, we have
@@
and the result follows.
We use Lemma 5.1 to show that our estimator of relative entropy differences has finite variance, at least under certain conditions on $\tilde{q}$.
Lemma B.2.
Assume that $\mathds{E}_{\mathds{P}}[\tau^{2}]<\infty$, that for some $C>0$,
[…]
@@
$\mathds{E}_{\mathds{P}}\left[\left(\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right)^{2}\right]<\infty$
even though $\frac{L\tilde{q}}{\tilde{q}}$ may be singular on $\partial A$. Taking $T=q$ in Lemma 5.1 yields
@@
where $r=\tilde{q}/q$ and $W_{s}$ is the $\mathds{P}$-Brownian motion in Theorem 3.11. Therefore,
$\mathds{E}_{\mathds{P}}\left[\left(\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right)^{2}\right]^{\frac{1}{2}}$ $\leq\mathds{E}_{\mathds{P}}\left[\left(\log\frac{r(Y_{\tau})}{r(Y_{0})}\right)^{2}\right]^{\frac{1}{2}}+\varepsilon\mathds{E}_{\mathds{P}}\left[\left(\int_{0}^{\tau}\lvert\nabla\log r(Y_{s})\rvert^{2}\,\mathrm{d}s\right)^{2}\right]^{\frac{1}{2}}$
$\qquad\qquad\qquad+\sqrt{2\varepsilon}\mathds{E}_{\mathds{P}}\left[\left(\int_{0}^{\tau}\nabla\log r(Y_{s})\cdot\mathrm{d}W_{s}\right)^{2}\right]^{\frac{1}{2}}$
(34) $\leq 2C+\varepsilon C\mathds{E}_{\mathds{P}}[\tau^{2}]+\sqrt{2\varepsilon}\mathds{E}_{\mathds{P}}\left[\left(\int_{0}^{\tau}\nabla\log r(Y_{s})\cdot\mathrm{d}W_{s}\right)^{2}\right]^{\frac{1}{2}}.$
@@
$\mathds{E}_{\mathds{P}}[(S_{\tau\wedge n^{\prime}}-S_{\tau\wedge n})^{2}]$ $=\mathds{E}_{\mathds{P}}\left[\left(\int_{\tau\wedge n}^{\tau\wedge n^{\prime}}\nabla\log r(Y_{s})\cdot\mathrm{d}W_{s}\right)^{2}\right]$
$=\mathds{E}_{\mathds{P}}\left[\left(\int_{0}^{n^{\prime}}\mathds{1}_{\tau\wedge n\leq s<\tau\wedge n^{\prime}}\nabla\log r(Y_{s})\cdot\mathrm{d}W_{s}\right)^{2}\right]$
$=\mathds{E}_{\mathds{P}}\left[\int_{0}^{n^{\prime}}\mathds{1}_{\tau\wedge n\leq s<\tau\wedge n^{\prime}}\lvert\nabla\log r(Y_{s})\rvert^{2}\,\mathrm{d}s\right]$
$\leq\mathds{E}_{\mathds{P}}\left[\int_{0}^{n^{\prime}}\mathds{1}_{\tau\geq n}\lvert\nabla\log r(Y_{s})\rvert^{2}\,\mathrm{d}s\right]$
@@
$\mathds{E}_{\mathds{P}}[S_{\tau}^{2}]$ $=\lim_{n\rightarrow\infty}\mathds{E}_{\mathds{P}}[S_{\tau\wedge n}^{2}]$
$=\lim_{n\rightarrow\infty}\mathds{E}_{\mathds{P}}\left[\left(\int_{0}^{n}\mathds{1}_{s\leq\tau\wedge n}\nabla\log r(Y_{s})\cdot\mathrm{d}W_{s}\right)^{2}\right]$
$=\lim_{n\rightarrow\infty}\mathds{E}_{\mathds{P}}\left[\int_{0}^{n}\mathds{1}_{s\leq\tau\wedge n}\lvert\nabla\log r(Y_{s})\rvert^{2}\,\mathrm{d}s\right]$
$\leq\lim_{n\rightarrow\infty}C^{2}\mathds{E}_{\mathds{P}}[\tau\wedge n]$
$=C^{2}\mathds{E}_{\mathds{P}}[\tau].$
@@
We leave the remaining details to the reader.
We now prove Theorem 4.5 on the differentiability of $D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ in $\theta$.
Proof B.4 (Proof of Theorem 4.5).
@@
Moreover, since we assume $\eta_{\theta}\geq\frac{1}{M}>0$, and since $q_{\theta}$ is differentiable, $\log\eta_{\theta}$ is differentiable. Therefore,
(35) $\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}$ $=\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\left(\frac{\nabla_{\theta}q_{\theta}(X_{\tau})}{q_{\theta}(X_{\tau})}-\frac{\nabla_{\theta}\eta_{\theta}}{\eta_{\theta}}-\int_{0}^{\tau}\nabla_{\theta}\frac{Lq_{\theta}}{q_{\theta}}(X_{s})\,\mathrm{d}s\right).$
@@
$D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})=\mathds{E}_{\mathds{Q}}\left[g\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)\right].$
@@
(36) $\nabla_{\theta}D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})=\nabla_{\theta}\mathds{E}_{\mathds{Q}}\left[g\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)\right]=\mathds{E}_{\mathds{Q}}\left[g^{\prime}\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right].$
Assumption 4.3 does not imply that $\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}$ is bounded uniformly in $\theta$ by a function in $L^{1}(\mathds{Q})$, so textbook results related to differentiating under the integral do not apply. To verify (36), we return to the definition of the derivative. Given any $h,\theta\in\mathbb{R}^{k}$, by the mean value theorem, for some $s:\Omega\rightarrow[0,1]$, we have
$\frac{1}{\lVert h\rVert}\mathds{E}_{\mathds{Q}}$ $\left[g\left(\frac{\mathrm{d}\mathds{P}_{\theta+h}}{\mathrm{d}\mathds{Q}}\right)-g\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)-g^{\prime}\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\cdot h\right]$
$=\mathds{E}_{\mathds{Q}}\left[g^{\prime}\left(\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\right)\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\cdot\frac{h}{\lVert h\rVert}-g^{\prime}\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\cdot\frac{h}{\lVert h\rVert}\right]$
$=\mathds{E}_{\mathds{Q}}\left[\left\{g^{\prime}\left(\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\right)-g^{\prime}\left(\frac{\mathrm{d}\mathds{P}_{\theta+h}}{\mathrm{d}\mathds{Q}}\right)\right\}\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\cdot\frac{h}{\lVert h\rVert}\right$.
@@
$=\mathds{E}_{\mathds{P}_{\theta}}\left[\left\{\log\left(\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\right)-\log\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)\right\}\nabla_{\theta}\log\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\cdot\frac{h}{\lVert h\rVert}\right$.
$\qquad\qquad+\left\{\log\left(\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\right)+1\right\}\nabla_{\theta}\log\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\cdot\frac{h}{\lVert h\rVert}\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\frac{\mathrm{d}\mathds{Q}}{\mathrm{d}\mathds{P}_{\theta}}$
@@
$=:\mathds{E}_{\mathds{P}_{\theta}}[R(\theta;h)].$
As $h\rightarrow 0$, the integrand $R(\theta;h)$ in the expectation with respect to $\mathds{P}_{\theta}$ above converges pointwise to zero. We now show that $R(\theta;h)$ is dominated in $L^{1}(\mathds{P}_{\theta})$, uniformly in $h$ for sufficiently small $h$. If so, $\lim_{h\rightarrow 0}\mathds{E}_{\mathds{P}_{\theta}}[R(\theta;h)]=0$, which verifies differentiability of $D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ and (36). By Assumption 4.3, we have
$\left\lvert\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\frac{\mathrm{d}\mathds{Q}}{\mathrm{d}\mathds{P}_{\theta}}\right\rvert$ $\leq\frac{\eta_{\theta}}{\eta_{\theta+sh}}\exp\left(\int_{0}^{\tau}\left\lvert\frac{Lq_{\theta}}{q_{\theta}}(Y_{t})-\frac{Lq_{\theta+sh}}{q_{\theta+sh}}(Y_{t})\right\rvert\,\mathrm{d}t\right)$
@@
[…]
$\leq\lvert\log\zeta\rvert+\lvert\log M\rvert+C\tau,$
@@
[…]
@@
$\lvert R(\theta;h)\rvert$ $\leq 2(\lvert\log\zeta\rvert+\lvert\log M\rvert+C\tau)(M^{2}+C\tau)$
$\qquad+(1+\lvert\log\zeta\rvert+\lvert\log M\rvert+C\tau)(M^{2}+C\tau)$
$\qquad+(1+\lvert\log\zeta\rvert+\lvert\log M\rvert+C\tau)(M^{2}+C\tau)M^{2}\exp\left(\frac{\gamma(\theta)}{2}\tau\right).$
Under Assumption 4.3, the random variable on the right-hand-side above is in $L^{1}(\mathds{P}_{\theta})$, which concludes the proof of differentiability.